Find the exact distance between the pair of points. (2, 0), and (4, -5)


1
Expert's answer
2022-01-25T14:47:49-0500

It follows that the exact distance between the pair of points "(2, 0)" and "(4, -5)" is equal to

"\sqrt{(4-2)^2+(-5-0)^2}=\sqrt{4+25}=\sqrt{29}."